# Multidata Causal Discovery for Statistical Hurricane Intensity Forecasting

### v2

@@ 3.3 Filtering new potential predictors
Figure 1 summarizes the causal feature selection and model evaluation framework. For all trained MLR models, we calculate the $R^{2}$ for training, validation, and test sets. For each cross-validation fold, we choose the model with the highest validation performance ($R^{2}$), resulting in 7 best performing models for analysis (upper-right panel in Figure 1). Comparing validation performance across all trained models allows us to identify the most generalizable MLR configuration. Adding more predictors beyond those chosen by the most generalizable MLR model will lead to model overfitting and make the MLR model less generalizable because they constitute spurious associations arising from autocorrelation, indirect effects, or common drivers. To ensure robustness across cases, the final causal variable list is obtained by aggregating and ranking the variables in the 7 best models. Only predictors that appear in at least four out of seven best models and are not part of the existing SHIPS developmental predictor list will be shortlisted as a candidate predictor (lower right panel in Figure 1).
@@ 3.4 Experimental Design
A hierarchy of experiments is conducted to demonstrate the potential applications of causal feature selection algorithms in complementing or enhancing existing statistical models for TC intensity change. We depict our overall experimental design in Figure 2.
@@ 3.4.1 ERA5-based experiments for predictor discovery
The first set of experiments involves identifying candidate causal predictors from ERA5 reanalysis data that can be incorporated into the operational SHIPS. We designed two complementary experiments to assess the algorithms and understand the sensitivity of discovered variables to the assumptions within the causal discovery process. In the first experiment (withASSUMPS), existing SHIPS predictors are retained a priori, and the causal discovery framework is restricted to identifying additional predictors. This represents a least-change and operationally feasible scenario. In contrast, the second experiment (noASSUMPS) is a free-run, “kitchen-sink” that allows the algorithm to freely evaluate all available predictors, including the possibility of excluding existing SHIPS predictors that do not exhibit strong causal relationships with TC intensity change. Through this experiment, we investigate whether removing spurious associations among the original SHIPS predictors improves the generalizability of statistical intensity prediction models. Causal feature selection is however constrained by the properties of the underlying datasets. ERA5 does not fully resolve the inner-core structure of tropical cyclones, and limitations in representing small-scale and azimuthally asymmetric processes are therefore reflected in the candidate predictors, particularly those related to the TC inner core. As a result, some physically relevant variables may not be identified as robust causal predictors within this framework.
@@ 3.5.3 Regression architecture
We first evaluate all predictor sets using multiple linear regression (MLR) models to preserve interpretability and isolate the impact of feature selection. For each experiment, we repeat 7-fold cross-validation across forecast lead times (24–120 hr), assessing performance on independent test sets. We further evaluate the predictor sets using multilayer perceptron (MLP) models, which we select based on their ability to capture nonlinear behaviors in the data that are not captured by MLR models. Each of the MLP models in our study includes a total of 5 layers, wherein each layer includes 512 units (also referred to as neurons). The first three layers rely on a Rectified Linear Unit (ReLU) activation function, the fourth layer instead uses hyperbolic tangent (tanh) to allow the model to output positive and negative intensity changes, and the final layer linearly combines the output of the 4th layer. All models were trained to minimize the mean square error (MSE) using the Adam optimizer Kingma and Ba (2014) with default parameters, a learning rate of 0.001, and with early stopping set to end training if the validation loss is greater than the average of the last 50 epochs once the model has trained for at least 50 epochs. This simplified architecture was chosen with rules of thumb and has not been subjected to a rigorous hyperparameter search given that the objective of these MLPs is not to achieve maximum performance, but rather to provide a simple nonlinear baseline for comparison to the MLRs.
@@ 4 Results
4.1 Effectiveness of multidata causal discovery for hurricane intensity prediction
Figure 3: Example results for the 24 hour intensity change forecast (DELV24) from Fold 3 using the SHIPS+ERA5 predictor set for SHIPS predictors. (a) Coefficient of determination $R^{2}$ on training, validation, and test sets plotted against the number of selected predictors, each point corresponding to a different value of the M-PC causal discovery hyperparameter $pc_{\alpha}$ (bottom scale). The vertical dashed line indicates the configuration with the highest validation $R^{2}$. (b) Variable selection abacus: each dot shows the presence of a predictor across the $pc_{\alpha}$ range in the test set. Variables are colored by group (e.g., Original SHIPS predictors, Shear, Humidity), vertical dashed line marks the best validation score, and encircled dots highlight the occurrence of new shortlisted predictors for SHIPS.
We now present results from the ERA5-based replication experiments, focusing on how the M-PC framework operates and how the predictor selections translate into statistical forecast skill using MLR. Figure 3 illustrates one realization of the ERA5-based causal discovery experiments for the 24 hour intensity change forecast (DELV24), shown for a representative cross-validation fold (Fold 3) under the noASSUMPS (“kitchen-sink”) configuration. The top panel shows training, validation, and test $R^{2}$ scores as a function of the number of input predictors in the MLR, controlled by the M-PC hyperparameter ${pc}_{\alpha}$. Higher values of ${pc}_{\alpha}$ correspond to a less strict statistical threshold for rejecting independence, allowing more predictors to be included in the causal model whereas a lower ${pc}_{\alpha}$ is more strict, resulting in a more stringent selection of predictors.
The bottom panel presents the variable selection abacus, highlighting predictor presence and groupings across the range of ${pc}_{\alpha}$ values. The encircled dots highlight the predictors included in the final shortlisted set, identified based on their consistent selection frequency across multiple folds. The same procedure was repeated for seven cross-validation folds, all forecast lead times, and both experimental setups. Given the volume of results, Figure 3 presents one representative fold for the noASSUMPS experiment, while the best performing folds for lead times of up to 120 hours are provided in the Supplementary Information (Figures S1–S10). The best potential causal predictors that complement or improve upon the baseline SHIPS predictors for TC intensity change forecasting tend to be consistently shortlisted by the causal feature selection framework across different data splits and validated configurations, ensuring that the predictors perform well over a wide range of TC cases.
@@ 4.2 Causal feature selection outperforms baseline methods
Figure 4: Summary of results for the 24 hour intensity change forecast (DELV24) using SHIPS+ERA5 predictors for the noASSUMPS experiment. (a) Bar plot showing the frequency of each variable’s selection across the best models from all seven cross-validation folds. A red dashed line marks the threshold (more than 3 folds) used to shortlist robust predictors for inclusion in the final SHIPS+ list. (b) Boxplot comparing test R² values for target DELV for each lead times 24, 48, 72, 96, 120 hrs for experiments with kitchen-sink approach (without link assumptions) across four feature selection strategies: causal discovery, correlation ranking, random forest importance, and no selection. Causal feature selection yields the highest median R² until 72 hrs lead time, showing improved generalization in a purely statistical prediction setup.
@@
Figure 4 summarizes the process and outcomes of identifying robust predictors using our causal feature selection pipeline for the 24 hour intensity change forecast (DELV24). Panel (a) shows how often each candidate predictor was selected in the seven cross-validation folds. By applying a clear threshold, which requires a predictor to appear in more than three folds—we filter out variables with inconsistent contributions and focus on predictors with stable links to intensity change. Panel (b) compares test $R^{2}$ scores for DELV24 across four feature selection strategies. Causal feature selection achieves the highest median skill, outperforming correlation ranking, random forest importance, and the no-selection baseline. This demonstrates that incorporating causally motivated predictors yields clear improvements in model generalizability. Comparing the causal and correlation feature selection methods, we observe that causal yields fewer outliers in prediction errors compared to the correlation-based approach, yet displays slightly larger uncertainty bounds. This behavior can be attributed to the fact that PC-based causal discovery method tends to identify stable, interventionally relevant relationships rather than purely predictive ones. When applied to a system with underlying nonlinear dynamics, the PC-based model may underfit complex regions of the predictor space, leading to broader prediction intervals, which reflects the model’s caution in extrapolating beyond what the causal structure supports. This caution mitigates against large, spurious prediction errors, and results in fewer outliers and more robust predictions on unseen test cases. The trade-off between model stability and uncertainty shown here highlights the value of causal methods over correlation for building reliable statistical models for high-impact forecasting applications. By extending this comparison across all lead times (Fig S11, S12 in the SI) and evaluating the consistency of variable selection, we ultimately shortlisted six predictors to be added to SHIPS. These variables reliably contribute to improved forecast skill for short-range intensity change prediction, particularly up to 72 hours. Beyond this time frame, the benefit of causal feature selection diminishes, emphasizing that while our method provides meaningful gains for in short-range, longer lead times will likely require integrating statistical models with dynamical forecast guidance to capture additional sources of predictability.
@@ 4.3 Recommending causally relevant predictors
In the previous section, we established that the causal predictors improve the generalizability of statistical TC intensity models for shorter lead times up to 72 hours but not for longer lead times (Fig. 4b).
@@
Before testing discovered variables in operational models, it is important to provide evidence that they are physically related to TC intensity change based on current knowledge. Low-tropospheric wind shear, which usually promotes linear convective organization rather than circular organization Robe and Emanuel (2001); Finocchio et al. (2016), is more negatively correlated with the intensity of Pacific typhoons that occurred in active typhoon seasons than commonly used deep-layer shear Wang et al. (2015).Low-level shear is further concluded to produce quasiperiodic oscillations in TC intensity that are related to variability in boundary-layer moisture induced by TC rainbands Fu et al. (2019). TC intensification is generally associated with relative humidity above the boundary layer Wu et al. (2012), as reflected by the inclusion of RHMD in the original SHIPS predictors. In contrast, the 1000 hPa relative humidity in the outer environment (R000) is a variable unused in operational models but available in the extended predictor list and selected by causal discovery. This selection can be physically explained as moisture within the boundary layer reflects the balance between surface flux moistening and dry air ventilation at the top of the boundary layer, through downdraft ventilation in sheared, dry environments Alland et al. (2021), which can strongly modulate convective buoyancy and intensification rates Riemer et al. (2010); Fischer et al. (2023). This shows that causal selection can reveal relevant predictors missed by traditional screening, improving short-range skills. This suggests that addition of overlooked factors like surface relative humidity alongside mid-tropospheric humidity and shear helps capture multiscale processes drive intensity change, supporting the idea that boundary layer moisture could play a key role in regulating TC intensities Wadler et al. (2021) and the onset of RI Chen et al. (2021). Finally, mid-tropospheric vortex strength is critical in the early intensification of TCs Bister and Emanuel (1997); Yang and Tan (2020).
@@ 4.4 Added value of the nonlinear SHIPS+ models
The results presented in previous sections suggested six additional predictors causally related to the change in TC intensity that need to be tested in the same setup as the operational SHIPS, where the GFS analysis and forecast fields are used to derive predictors. As a first step towards a fair evaluation of operational model skill, we compare models trained with the original SHIPS predictors and the enriched SHIPS+, which includes the six newly identified causal predictors. For consistency with the ERA5 experiments, only the initial analysis time data are used to derive both SHIPS developmental and SHIPS+ datasets. This differs from the operational SHIPS setup, which incorporates dynamical model forecast output, so the results here may be negatively biased at longer lead times when synoptic-scale environmental changes are important.
@@
Figure 5: Comparison of test $R^{2}$ values across forecast lead times (24–120 hr) for the original SHIPS predictors (blue/green boxes) and the expanded SHIPS+ predictors (orange/yellow boxes), using no feature selection. Both MLR and MLP runs are shown to illustrate the added value of nonlinear modeling. Dashed brown lines indicate the median, and solid black lines mark the mean. Overall, the MLP consistently outperforms the MLR, demonstrating improved skill when nonlinearity is captured, while the inclusion of additional predictors in SHIPS+ further enhances forecast performance, especially at shorter lead times.Note that the SHIPS+ MLR $R^{2}$ drops below 0 at the 120 hr lead time, which is indicated in the figure by a downward arrow.
@@
Using $R^{2}$ as the performance metric, regression models trained with the additional causal predictors (yellow and orange boxes in Fig. 5) outperform models trained with only the original SHIPS predictors (green and blue boxes in Fig. 5) on unseen test TC cases, particularly at shorter lead times. This improvement is seen in both the linear MLR (blue and orange boxes) and nonlinear MLP (green and yellow boxes) models. However, the nonlinear SHIPS+ MLPs generally outperform their linear MLR counterparts. Since the SHIPS model is fully linear, we are interested in determining whether the superior MLP test skills can instead be attributed to the inability of MLRs to capture nonlinear relationships between the causal predictors and TC intensity changes. Indeed, the lead-time dependence of test skill improvements changes significantly in a nonlinear regression framework. The SHIPS+ MLRs show a clear lead-time dependence in the effect of the additional causal predictors, performing worse than the original SHIPS model beyond 72 hours, with negative $R^{2}$ at 120 hours. This lead time dependence is reduced in a nonlinear regression framework, where SHIPS+ outperforms SHIPS for all lead times. The results in Figure 5 implies that the relationship between the identified causal predictors and the rate of intensity changes of the TC can be approximated linearly at shorter lead times but not at extended lead times, which makes the MLRs incapable of utilizing the additional information beyond 72 hours. Finally, the wider uncertainty ranges of the trained models at longer lead times suggest that long-lead-time predictions cannot be adequately constrained with predictors derived solely from analysis time. This is a fundamental limitation of our predictive modeling setup, but it could be alleviated in operational settings by using dynamical model forecasts.
@@
To better understand why the nonlinear SHIPS+ models outperform the other models, we conducted a SHAP (SHapley Additive exPlanations) analysis for the 24 hour lead time MLP and MLR models using both SHIPS and SHIPS+ datasets (see Fig. 6). SHAP values provide a model-agnostic interpretation of feature importance by quantifying the marginal contribution of each predictor to the model’s output. To conduct the SHAP analysis, we rely on Kernel SHAP Lundberg and Lee (2017), using 300 samples from the training dataset to retrieve the background signal and evaluating on the 240 samples in the test set. Focusing on the differences between the linear MLR and the MLP models, we observe that of the six causal predictors, the MLP primarily utilizes the near-surface moisture predictors (red bars in Fig.6a), whereas the SHIPS+ MLRs (Fig.6b) use the causal shear predictors instead. A potential reason for MLPs and MLRs to select different causal predictors is that the relationship between near-surface moisture and short-term changes in TC intensity is too nonlinear to be used by linear MLRs. The dependencies of the causal boundary layer moisture predictors (e.g., R001) on the MLR predictions are nonlinear (Fig. 6e). Interestingly, the SHAP dependence plots reveal that the MLP learns opposing relationships for inner-core and outer-core near-surface humidity (Fig. 6e): increases in R001 (0–500 km) are associated with weaker predicted intensification, whereas increases in R000 (200–800 km) support stronger intensification. This contrast indicates that the MLP is leveraging radial gradients in boundary-layer moisture, a signal that the linear MLR framework does not capture. Apart from prioritizing additional nonlinear dependencies, the SHIPS+ MLPs could also overperform their MLR counterparts by learning different dependencies for the existing SHIPS predictors. Figure 6d presents two examples in which this is the case, where low-level wind shear (SHL1) and mid-tropospheric potential vorticity (PVOR) have a much stronger dependence on MLP predictions than the muted ones for MLR models. The negative correlation between intensification and PVOR might reflect the negative impact of TC outer rainbands on intensification Wang (2009); Yu et al. (2021). Furthermore, MLPs learned a dependence between SHL1 and short-term intensity evolution that is strikingly different from the MLR dependence. Although intensification predictions in both the nonlinear and linear regression frameworks have a similar dependency on potential intensity (POT), which is the most critical predictor in both types of models, the difference in the selection of other leading predictors and their respective learned dependencies contribute to the significant differences in the generalizability of the model to test TCs (Fig.  5).

### v3

@@ 3.3 Filtering new potential predictors
Fig.  1 summarizes the causal feature selection and model evaluation framework. For all trained MLR models, we calculate the $R^{2}$ for training, validation, and test sets. For each cross-validation fold, we choose the model with the highest validation performance ($R^{2}$), resulting in 7 best performing models for analysis (upper-right panel in Fig. 1). Comparing validation performance across all trained models allows us to identify the most generalizable MLR configuration. Adding more predictors beyond those chosen by the most generalizable MLR model will lead to model overfitting and make the MLR model less generalizable because they constitute spurious associations arising from autocorrelation, indirect effects, or common drivers. To ensure robustness across cases, the final causal variable list is obtained by aggregating and ranking the variables in the 7 best models. Only predictors that appear in at least four out of seven best models and are not part of the existing SHIPS developmental predictor list will be shortlisted as a candidate predictor (lower right panel in Fig. 1).
@@ 3.4 Experimental Design
A hierarchy of experiments is conducted to demonstrate the potential applications of causal feature selection algorithms in complementing or enhancing existing statistical models for TC intensity change. We depict our overall experimental design in Fig.  2.
@@ 3.4.1 ERA5-based experiments for predictor discovery
The first set of experiments involves identifying candidate causal predictors from ERA5 reanalysis data that can be incorporated into the operational SHIPS. We designed two complementary experiments to assess the algorithms and understand the sensitivity of discovered variables to the assumptions within the causal discovery process. In the first experiment (withASSUMPS), existing SHIPS predictors are retained a priori, and the causal discovery framework is restricted to identifying additional predictors. This represents a least-change and operationally feasible scenario. In contrast, the second experiment (noASSUMPS) is a free-run, “kitchen-sink” that allows the algorithm to freely evaluate all available predictors, including the possibility of excluding existing SHIPS predictors that do not exhibit strong causal relationships with TC intensity change. Through this experiment, we investigate whether removing spurious associations among the original SHIPS predictors improves the generalizability of statistical intensity prediction models. Causal feature selection is however constrained by the properties of the underlying datasets. ERA5 does not fully resolve the inner-core structure of TCs, and limitations in representing small-scale and azimuthally asymmetric processes are therefore reflected in the candidate predictors, particularly those related to the TC inner core. As a result, some physically relevant variables may not be identified as robust causal predictors within this framework.
@@ 3.5.3 Regression architecture
We first evaluate all predictor sets using multiple linear regression (MLR) models to preserve interpretability and isolate the impact of feature selection. For each experiment, we repeat 7-fold cross-validation across forecast lead times (24–120 hours), assessing performance on independent test sets. We further evaluate the predictor sets using multilayer perceptron (MLP) models, which we select based on their ability to capture nonlinear behaviors in the data that are not captured by MLR models. Each of the MLP models in our study includes a total of 5 layers, wherein each layer includes 512 units (also referred to as neurons). The first three layers rely on a Rectified Linear Unit (ReLU) activation function, the fourth layer instead uses hyperbolic tangent (tanh) to allow the model to output positive and negative intensity changes, and the final layer linearly combines the output of the 4th layer. All models were trained to minimize the mean square error (MSE) using the Adam optimizer Kingma and Ba (2014) with default parameters, a learning rate of 0.001, and with early stopping set to end training if the validation loss is greater than the average of the last 50 epochs once the model has trained for at least 50 epochs. This simplified architecture was chosen with rules of thumb and has not been subjected to a rigorous hyperparameter search given that the objective of these MLPs is not to achieve maximum performance, but rather to provide a simple nonlinear baseline for comparison to the MLRs.
@@ 4 Results
4.1 Effectiveness of multidata causal discovery for TC intensity prediction
Figure 3: Example results for the 24-hour intensity change forecast (DELV24) from Fold 3 using the SHIPS+ERA5 predictor set for SHIPS predictors. (a) Coefficient of determination $R^{2}$ on training, validation, and test sets plotted against the number of selected predictors, each point corresponding to a different value of the M-PC causal discovery hyperparameter $pc_{\alpha}$ (bottom scale). The vertical dashed line indicates the configuration with the highest validation $R^{2}$. (b) Variable selection abacus: each dot shows the presence of a predictor across the $pc_{\alpha}$ range in the test set. Variables are colored by group (e.g., Original SHIPS predictors, Shear, Humidity), vertical dashed line marks the best validation score, and encircled dots highlight the occurrence of new shortlisted predictors for SHIPS.
We now present results from the ERA5-based replication experiments, focusing on how the M-PC framework operates and how the predictor selections translate into statistical forecast skill using MLR. Fig.  3 illustrates one realization of the ERA5-based causal discovery experiments for the 24-hour intensity change forecast (DELV24), shown for a representative cross-validation fold (Fold 3) under the noASSUMPS (“kitchen-sink”) configuration. The top panel shows training, validation, and test $R^{2}$ scores as a function of the number of input predictors in the MLR, controlled by the M-PC hyperparameter ${pc}_{\alpha}$. Higher values of ${pc}_{\alpha}$ correspond to a less strict statistical threshold for rejecting independence, allowing more predictors to be included in the causal model whereas a lower ${pc}_{\alpha}$ is more strict, resulting in a more stringent selection of predictors.
The bottom panel presents the variable selection abacus, highlighting predictor presence and groupings across the range of ${pc}_{\alpha}$ values. The encircled dots highlight the predictors included in the final shortlisted set, identified based on their consistent selection frequency across multiple folds. The same procedure was repeated for seven cross-validation folds, all forecast lead times, and both experimental setups. Given the volume of results, Fig.  3 presents one representative fold for the noASSUMPS experiment, while the best performing folds for lead times of up to 120 hours are provided in the Supplementary Information (Figures S1–S10). The best potential causal predictors that complement or improve upon the baseline SHIPS predictors for TC intensity change forecasting tend to be consistently shortlisted by the causal feature selection framework across different data splits and validated configurations, ensuring that the predictors perform well over a wide range of TC cases.
@@ 4.2 Causal feature selection outperforms baseline methods
Figure 4: Summary of results for the 24-hour intensity change forecast (DELV24) using SHIPS+ERA5 predictors for the noASSUMPS experiment. (a) Bar plot showing the frequency of each variable’s selection across the best models from all seven cross-validation folds. A red dashed line marks the threshold (more than 3 folds) used to shortlist robust predictors for inclusion in the final SHIPS+ list. (b) Boxplot comparing test R² values for target DELV for each lead time (24, 48, 72, 96, 120 hours) for experiments with kitchen-sink approach (without link assumptions) across four feature selection strategies: causal discovery, correlation ranking, random forest importance, and no feature selection (all features). Causal feature selection yields the highest median R² until 72-hour lead time, showing improved generalization in a purely statistical prediction setup.
@@
Fig.  4 summarizes the process and outcomes of identifying robust predictors using our causal feature selection pipeline for the 24-hour intensity change forecast (DELV24). Panel (a) shows how often each candidate predictor was selected in the seven cross-validation folds. By applying a clear threshold, which requires a predictor to appear in more than three folds, we filter out variables with inconsistent contributions and focus on predictors with stable links to intensity change. Panel (b) compares test $R^{2}$ scores for DELV across the four feature selection strategies at different lead times for the noASSUMPS experiment. Corresponding train and validation results are shown in Fig. S14, while results for the withASSUMPS experiment (train, validation, and test) are provided in Fig. S13 of the Supplementary Information. Causal feature selection achieves the highest median skill, outperforming correlation ranking, random forest importance, and the no feature selection baseline, demonstrating improved generalizability with causal predictors. Comparing the causal and correlation feature selection methods, we observe that causal yields fewer outliers in prediction errors compared to the correlation-based approach, yet displays slightly larger uncertainty bounds. This is because the PC-based causal discovery method tends to identify stable, interventionally relevant relationships rather than purely predictive ones. When applied to a system with underlying nonlinear dynamics, the PC-based model may underfit complex regions of the predictor space, leading to broader prediction intervals, which reflects the model’s caution in extrapolating beyond what the causal structure supports. This caution mitigates large, spurious prediction errors and results in fewer outliers and more robust predictions on unseen test cases. The trade-off between model stability and uncertainty shown here highlights the value of causal methods over correlation in building reliable statistical models for high-impact forecasting applications. By extending this comparison across all lead times (Fig S11, S12 in the SI) and evaluating the consistency of variable selection, we ultimately shortlisted six predictors to be added to SHIPS. These variables reliably contribute to improved forecast skill for short-range intensity change prediction, particularly up to 72 hours. Beyond this time frame, the benefit of causal feature selection diminishes, emphasizing that while our method provides meaningful gains in the short-range, longer lead times will likely require integrating statistical models with dynamical forecast guidance to capture additional sources of predictability.
@@ 4.3 Recommending causally relevant predictors
In the previous section, we established that the causal predictors improve the generalizability of statistical TC intensity models for shorter lead times up to 72 hours, but not for longer lead times (Fig. 4b).
@@
The physical relevance of these variables to TC intensity change is assessed prior to operational testing. Low-tropospheric wind shear, which usually promotes linear convective organization rather than circular organization Robe and Emanuel (2001); Finocchio et al. (2016), is more negatively correlated with the intensity of Pacific typhoons that occurred in active typhoon seasons than commonly used deep-layer shear Wang et al. (2015). (Fu et al., 2019) further conclude that low-level shear produces quasiperiodic oscillations in the intensity of the TC that are related to the variability in the moisture of the boundary layer induced by the TC rainbands. TC intensification is generally associated with relative humidity above the boundary layer Wu et al. (2012), as reflected by the inclusion of RHMD in the original SHIPS predictors. In contrast, the 1000 hPa relative humidity in the outer environment (R000) is a variable not used in operational models but available in the extended predictor list. This selection can be physically explained as moisture within the boundary layer reflects the balance between surface flux moistening and dry air ventilation at the top of the boundary layer, through downdraft ventilation in sheared, dry environments Alland et al. (2021), which can strongly modulate convective buoyancy and intensification rates Riemer et al. (2010); Fischer et al. (2023). Causal selection therefore reveals relevant predictors missed by traditional screening, improving short-range skills. The addition of overlooked variables such as surface relative humidity along with mid-tropospheric humidity and shear helps capture multiscale processes that drive intensity change, supporting the idea that boundary layer moisture could play a key role in regulating TC intensities Wadler et al. (2021) and the onset of RI Chen et al. (2021). Mid-level potential vorticity in the outer region of TCs reflects the dynamical environment of the storm and may capture interactions with environmental PV anomalies known to influence TC intensity Molinari et al. (1998), including synoptic-scale influences such as trough interactions Fischer et al. (2019).
@@ 4.4 Added value of the nonlinear SHIPS+ models
The previous sections identify six additional predictors causally related to TC intensity change. These are evaluated within the operational SHIPS framework using predictors derived from GFS analysis and forecast fields. As a first step towards a fair evaluation of operational model skill, we compare models trained with the original SHIPS predictors and the enriched SHIPS+, which includes the six predictors. For consistency with the ERA5 experiments, only the initial analysis time data are used to derive both SHIPS developmental and SHIPS+ datasets.
@@
Figure 5: Comparison of test $R^{2}$ values across forecast lead times (24–120 hours) for the original SHIPS predictors (blue/green boxes) and the expanded SHIPS+ predictors (orange/yellow boxes). Both MLR and MLP runs are shown to illustrate the added value of nonlinear modeling. Dashed brown lines indicate the median, and solid black lines mark the mean. Overall, the MLP consistently outperforms the MLR, demonstrating improved skill when nonlinearity is captured, while the inclusion of additional predictors in SHIPS+ further enhances forecast performance, especially at shorter lead times. Note that the SHIPS+ MLR $R^{2}$ drops below 0 at the 120-hour lead time, which is indicated in the figure by a downward arrow.
@@
Using $R^{2}$ as the performance metric, regression models trained with the additional causal predictors (yellow and orange boxes in Fig.  5) outperform models trained with only the original SHIPS predictors (green and blue boxes in Fig.  5) on unseen test TC cases, particularly at shorter lead times. This improvement is seen in both the linear MLR (blue and orange boxes) and nonlinear MLP (green and yellow boxes) models. However, the nonlinear SHIPS+ MLPs generally outperform their linear MLR counterparts. Since the SHIPS model is fully linear, we assess whether the superior performance of MLP models arises from their ability to capture nonlinear relationships between causal predictors and TC intensity change. Indeed, the lead time dependence of test skill improvements changes significantly in a nonlinear regression framework. The SHIPS+ MLRs show a clear lead time dependence with additional causal predictors, performing worse than the original SHIPS model beyond 72 hours, with negative $R^{2}$ at 120 hours. This lead time dependence is reduced in the nonlinear regression framework, where SHIPS+ outperforms SHIPS for all lead times. The results in Figure 5 imply that the relationship between the identified causal predictors and the rate of TC intensity change can be approximated linearly at shorter lead times but not at extended lead times, which makes MLRs incapable of utilizing additional information beyond 72 hours. Finally, the wider uncertainty ranges of the trained models at longer lead times suggest that these predictions cannot be adequately constrained with predictors derived solely from analysis time. This is a fundamental limitation of our predictive modeling setup, but it could be alleviated in operational settings by using dynamical model forecasts.
@@
To better understand why the nonlinear SHIPS+ models outperform the other models, we conducted a SHAP (SHapley Additive exPlanations) analysis for the 24-hour lead time MLP and MLR models using both SHIPS and SHIPS+ datasets (see Fig.  6). SHAP values provide a model-agnostic interpretation of feature importance by quantifying the marginal contribution of each predictor to the model’s output. To perform the SHAP analysis, we rely on Kernel SHAP Lundberg and Lee (2017), using 300 samples from the training dataset to retrieve the background signal and evaluating the 240 samples in the test set. Focusing on the differences between the linear MLR and the MLP models, we observe that of the six causal predictors, the MLP primarily utilizes the near-surface moisture predictors (red bars in Fig. 6a), while the SHIPS+ MLRs (Fig. 6b) use the causal shear predictors instead. A potential reason for MLPs and MLRs to select different causal predictors is that the relationship between near-surface moisture and short-term changes in TC intensity is too nonlinear for linear MLRs. The dependencies of the causal boundary layer moisture predictors (e.g., R001) on the MLR predictions are nonlinear (Fig. 6e). Interestingly, the SHAP dependence plots reveal that the MLP learns opposing relationships for inner-core and outer-core near-surface humidity (Fig.  6e): increases in R001 (0–500 km) are associated with weaker predicted intensification, whereas increases in R000 (200–800 km) support stronger intensification. This result is unexpected, as higher inner-region moisture is generally associated with intensification. This contrast indicates that the MLP is leveraging radial gradients in boundary-layer moisture, a signal that the linear MLR framework does not capture. Apart from prioritizing additional nonlinear dependencies, the SHIPS+ MLPs could also overperform their MLR counterparts by learning different dependencies for the existing SHIPS predictors. Figure 6d presents two examples in which this is the case, where low-level wind shear (SHL1) and mid-tropospheric potential vorticity (PVOR) have a much stronger dependence on MLP predictions than the muted ones for MLR models. The negative correlation between intensification and PVOR might reflect the negative impact of TC outer rainbands on intensification Wang (2009); Yu et al. (2021). Furthermore, MLPs learned a dependence between SHL1 and short-term intensity evolution that is strikingly different from the MLR dependence. Although intensification predictions in both the nonlinear and linear regression frameworks have a similar dependency on potential intensity (POT), which is the most critical predictor in both types of models, the difference in the selection of other leading predictors and their respective learned dependencies contribute to significant differences in the generalizability of the model to test TCs (Fig.  5).